# Closed Orbits of Dynamically Convex Reeb Flows: Towards the HZ- and Multiplicity Conjectures

### v2

@@ 4.3.3. Proof of Corollary D
The first assertion is a consequence of (2.11), (2.21) and Theorem C. Indeed, if we had two ${\mathbb{F}}$-visible orbits $x$ and $x^{\prime}$ with action $a$, by (2.11) and (2.21), we would have at least four bars with end-point $a$ and hence $\dim\operatorname{SH}^{t}(W)\geq 2$ for some $t$ close to $a$. Now the second assertion follows immediately from (2.10) and the first one. ∎
@@ 4.4. Proof of Corollary F
As was already mentioned in Section 1.3, by Part (O2) of Theorem E, we can scale $M$ so that $\operatorname{\hat{\mu}}(x)={\mathcal{A}}(x)$ for all closed orbits in $M$. From now on we assume that this condition is satisfied. Denote by $x_{1},\ldots,x_{n}$ the prime closed orbits in $M$. Then the action spectrum of $M$ is the same as the mean index spectrum and is simply the union of the sequences $\Delta_{j}{\mathbb{N}}$. Note also that we have $\Delta_{j}/\Delta_{i}\not\in{\mathbb{Q}}$, when $i\neq j$, by again (O2), since all closed orbits have distinct actions due to Corollary D.
@@
Recall that the non-resonance condition is that for every $j$ the ratios ±$\pm\Delta_{j}/\Delta_{i}=\pm\lambda_{ij}$ are distinct modulo ${\mathbb{Z}}$ for every fixed $j$ and all $i\neq j$. This guarantees that the first Krein type eigenvalues are distinct for every $j$ and are not complex conjugate to each other.
@@
By [Vi89], the sum in this formula is $1/2$; see also [GiK]. Therefore,
@@
Without the non-resonance condition, the argument still provides some comparison information about the linearized flows along $x_{j}$ and $y_{j}$. Namely, it shows that the return maps have the same first Krein type eigenvalues, counted with multiplicity which is determined as follows. Assume that a linear symplectic map has both $w$ and $\bar{w}$ as first Krein type eigenvalues and their algebraic multiplicities are $m$ and, respectively, $l$. By definition, we set the multiplicity of $w$ to be $m-l$ and the multiplicity of $\bar{w}$ to be $l-m$; see [GG22, Sec. 4]. (Note that with this way of counting multiplicity an elliptic orbit such that its first Krein type eigenvalues come in complex conjugate pairs is indistinguishable from a hyperbolic orbit.) Furthermore, as in the non-resonance case, $x_{j}$ and $y_{j}$ have the same mean index (see (4.16)) and also, up to a factor, the same action.
@@ 5.1. Input from geometry of numbers
Let $U\subset{\mathbb{T}}^{n}={\mathbb{R}}^{n}/{\mathbb{Z}}^{n}$ be a neighborhood of 00 and $\theta\in{\mathbb{T}}^{n}$. Consider the sequence $k_{1}\leq k_{2}\leq k_{3}\leq\ldots$ of all positive integers such that $k_{l}\theta\in U$. Clearly, $k_{l}\to\infty$. Moreover, it is well known that the orbit $\{k\theta\mid k\in{\mathbb{N}}\}$ is equidistributed in the closure $G$ of $\theta{\mathbb{Z}}$ in ${\mathbb{T}}^{n}$. In particular, the density of this sequence in ${\mathbb{N}}$ is equal to the $\dim G$-dimensional volume of $U\cap G$.
@@
The strategy of the proof is as follows. We first show in Section 5.2.2 that for any element $\Phi:=\Phi_{ij}$ assertions (IR1)(IR5) are formal consequences of the requirement that $k=k_{ijl}$ satisfies a certain system of inequalities completely determined by $\Phi$. Here $d:=d_{il}$ is determined by $k$ and, of course, $\Phi$ by the first part of (IR1). Furthermore, (IR5) can be taken as the defining condition for $C_{l}$. (The existence of a bounded gap infinite sequence of solutions $k$ of this system of inequalities is a consequence of Lemma 5.1.)
@@
5.2.1. General observations
@@
The claim that the sequences $C_{l}$, $d_{il}$ and $k_{ijl}$ can be taken strictly monotone increasing is a formal consequence of the rest of the theorem. Indeed, observe that every bounded gap sequence $d_{il}\to\infty$ has a bounded gap strictly monotone increasing refinement $s\mapsto d_{il_{s}}$. Moreover, for any $c>0$ we may ensure by further thinning the sequence that $d_{il_{s+1}}-d_{l_{s}}>c$. Hence, by passing to subsequences of $d_{il}$ and $k_{ijl}$, we can guarantee that $d_{il}$ is strictly monotone increasing, $d_{il+1}-d_{il}>c$ and these sequences still have bounded gap. Then, since $\operatorname{\hat{\mu}}\big{(}\Phi^{k_{ijl}}_{ij}\big{)}=k_{ijl}\operatorname{\hat{\mu}}(\Phi_{ij})$ and $\operatorname{\hat{\mu}}(\Phi_{ij})>0$, we conclude from (IR1) that the sequences $k_{ijl}$ are automatically strictly monotone increasing when $c$ is sufficiently large. Likewise, the sequence $C_{l}$ is forced to be strictly monotone increasing by (IR5) once $\eta>0$ is sufficiently small.
@@
Let us decompose $\Phi$ as the product of a loop $\varphi$ and a direct sum $\Phi_{\scriptscriptstyle{deg}}\oplus\Phi_{h}\oplus\Phi_{e}$ with the following properties. The end-point $\Phi_{\scriptscriptstyle{deg}}(1)$ is totally degenerate and $\operatorname{\hat{\mu}}(\Phi_{\scriptscriptstyle{deg}})=0$. The path $\Phi_{h}$ has a hyperbolic end-point $\Phi_{h}(1)$, while $\Phi_{e}$ is the elliptic part of $\Phi$, i.e., all eigenvalues of $\Phi_{e}(1)$ are on the unit circle and different from 1 due to the presence of $\Phi_{\scriptscriptstyle{deg}}$. We write the eigenvalues of $\Phi_{e}(1)$ as $\exp(\pm 2\pi\sqrt{-1}\lambda_{q})$, where $20<|\lambda_{q}|\leq 1/$ 2 and $q=1,\ldots,m^{\prime}$. Here for each pair of complex conjugate eigenvalues it is convenient to pick the sign of $\lambda_{q}$ so that the resulting eigenvalue is of the first Krein type; see [SZ]. (We will elaborate on this choice later.) Furthermore, we then require that
@@
[…]
@@
Before proving the lemma, we note that (5.5) has a bounded gap sequence of solutions $k_{l}\to\infty$ by Lemma 5.1 and then, as we have already pointed out, the corresponding sequence $d=d_{l}\to\infty$ has bounded gap automatically. For any $a:=a_{ij}>0$ to find a bounded gap sequence $C_{l}\to\infty$ meeting condition (IR5), we can just pick any $C_{l}\not\in a{\mathbb{N}}$ in the range $(k_{l}a-\eta,k_{l}a)$ assuming that $\eta<a$. Furthermore, to make sure that $k$ is divisible by $D$ or any fixed integer $N>0$, it suffices to replace $\epsilon$ by $\epsilon/N$ in the definition of $U$. Then, after applying Lemma 5.1 for this new neighborhood $U$, we replace sequence of solutions $k$ by $Nk$ and have a bounded gap sequence of solutions satisfying (5.5) and divisible by $N$. As the next step, to meet the requirement of (5.6) we have $d$ replaced by $Nd$, and hence also divisible by $N$, provided that $\epsilon>0$ is sufficiently small. Thus, to complete this step of the proof, it remains to prove Lemma 5.3. The proof has a substantial overlap with [GG20, Sec. 5.2.1] and is implicitly contained in that section.
@@
[…]
@@
We will use this observation to reduce the proof to several particular cases. Throughout the argument we will assume without specifically stating this that $k\in{\mathbb{N}}$ and $\ell\in{\mathbb{N}}$ and $k>\ell$ in (IR3”).
@@
Next, observe that without loss of generality we may assume that map $\Phi(1)$ is semisimple. Indeed, consider an isospectral (i.e., with constant eigenvalues) family $\Psi_{s}$, $s\in[0,1]$, connecting $\Phi(1)$ to a semisimple linear symplectic map with the same eigenvalues as $\Phi(1)$. Let $\tilde{\Phi}_{s}\in\widetilde{\operatorname{Sp}}(2m)$ be obtained by concatenating $\Phi$ and $\Psi$. Then $\mu\big{(}\tilde{\Phi}_{s}^{k}\big{)}=\mu\big{(}\Phi^{k}\big{)}$ and $\operatorname{\hat{\mu}}\big{(}\tilde{\Phi}_{s}^{k}\big{)}=\operatorname{\hat{\mu}}\big{(}\Phi_{s}^{k}\big{)}$ for all $k$ and $s$, and hence in the proof of the theorem we can replace $\Phi$ by $\tilde{\Phi}$. We keep the notation $\Phi$.
@@
[…]
@@
[…]
@@
[…]
@@
In spite of its appearance this statement is not entirely obvious and requires a proof. An easy way to prove (5.13) is as follows. Pick an isospectral family $\Psi_{s}$, $s\in[0,1]$, of transformations with $\Psi_{1}=\Phi_{1}$ and $\Psi_{0}(1)$ semisimple. Then $\Psi_{0}^{k}(1)={\mathit{i}d}$, i.e., $\Psi_{0}^{k}$ is a loop, and $\operatorname{\hat{\mu}}\big{(}\Phi_{1}^{k}\big{)}=\operatorname{\hat{\mu}}\big{(}\Psi_{0}^{k}\big{)}$ because all eigenvalues of $\Psi_{s}^{k}(1)$ are equal to 1 for all $s$. Furthermore, $\Psi_{s}^{k\pm\ell}$ and $\Psi_{s}^{\ell}$ remain non-degenerate for all $s$. As a consequence, $\mu\big{(}\Phi_{1}^{\ell}\big{)}=\mu\big{(}\Psi_{0}^{\ell}\big{)}$ and $\mu\big{(}\Phi_{1}^{k\pm\ell}\big{)}=\mu\big{(}\Psi_{0}^{k\pm\ell}\big{)}$. It follows that (5.13) holds if and only if it holds for $\Psi_{0}$. The latter, however, is obvious:
@@
Recall also that
$\beta_{\pm}\big{(}\Phi_{0}^{k+\ell}\big{)}=\beta_{\pm}(\Phi_{0})=\beta_{\pm}\big{(}\Phi_{0}^{\ell}\big{)}$
by (3.4). Therefore, using (5.13) in the third line below, we have
@@
Then a simple examination of the rest of the proof shows that (IR1”)(IR3”) still hold for $\Psi$. ∎
@@
As in Section 5.2.2, denote the elliptic eigenvalues of $\Phi_{ij}$ by $\exp(\pm 2\pi\sqrt{-1}\lambda_{ijq})$, where $20<|\lambda_{ijq}|\leq 1/$ 2 and $q=1,\ldots,m_{ij}^{\prime}$, and for each pair of complex conjugate eigenvalues we pick the sign of $\lambda_{ijq}$ so that the resulting eigenvalue is of the first Krein type; see [SZ]. Set $\Delta_{ij}:=\operatorname{\hat{\mu}}(\Phi_{ij})$ and let $D$ be the least common multiple of the degrees of the roots of unity among the eigenvalues of all $\Phi_{ij}$.
@@
$|k_{11}a_{11}-k_{ij}a_{ij}|$ $<\sigma\quad\textrm{ for $i\geq 2$ and all $j$,}$

### v3

@@ 4.3.3. Proof of Corollary D
The first assertion is a consequence of (2.12), (2.24) and Theorem C. Indeed, if we had two $\mathbb{F}$}-visible orbits $x$ and $x^{\prime}$ with action $a$, by (2.12) and (2.24), we would have at least four bars with end-point $a$ and hence $\dim\operatorname{SH}^{t}(W)\geq 2$ for some $t$ close to $a$. Likewise, $\dim\operatorname{SH}(x;{\mathbb{F}})=2$ for all $\mathbb{F}$}-visible orbits by (2.24) and Theorem C. Now the second assertion follows immediately from (2.10) and the first one. ∎
@@ 4.4. Proof of Corollary F
As was already mentioned in Section 1.3, by Part (O2) of Theorem E, we can scale $M$ so that $\operatorname{\hat{\mu}}(x)={\mathcal{A}}(x)$ for all closed orbits in $M$. From now on we assume that this condition is satisfied. Denote by $x_{1},\ldots,x_{n}$ the prime closed orbits in $M$. Then the action spectrum of $M$ is the same as the mean index spectrum and is simply the union of the sequences $\Delta_{j}{\mathbb{N}}$. Note also that we have $\Delta_{j}/\Delta_{i}\not\in{\mathbb{Q}}$, when $i\neq j$, again by (O2), since all closed orbits have distinct actions due to Corollary D.
@@
Recall that the non-resonance condition is that for every $j$ the ratios $\pm\Delta_{j}/\Delta_{i}=\pm\lambda_{ij}$ are distinct modulo $\mathbb{Z}$} for every fixed $j$ and all $i\neq j$. This guarantees that the first Krein type eigenvalues are distinct for every $j$ and are not complex conjugate to each other.
@@
As a consequence of Theorem A, the prime orbits $x_{1},\ldots,x_{n}$ are non-alternating. By [Vi89], the sum in this formula is $1/2$; see also [GiK]. Therefore,
@@
Without the non-resonance condition, the argument still provides some comparison information about the linearized flows along $x_{j}$ and $y_{j}$. Namely, it shows that the return maps have the same first Krein type eigenvalues, counted with multiplicity which is defined as follows. Assume that a linear symplectic map has both $w$ and $\bar{w}$ as first Krein type eigenvalues and their algebraic multiplicities are $m$ and, respectively, $l$. Then we set, by definition, the multiplicity of $w$ to be mlm-l and the multiplicity of $\bar{w}$ to be lml-m; see [GG22, Sec. 4]. (Note that with this way of counting multiplicity an elliptic orbit such that its first Krein type eigenvalues come in complex conjugate pairs is indistinguishable from a hyperbolic orbit.) Furthermore, as in the non-resonance case, $x_{j}$ and $y_{j}$ have the same mean index (see (4.17)) and also, up to a factor, the same action.
@@ 5.1. Input from geometry of numbers
Let $U\subset{\mathbb{T}}^{n}={\mathbb{R}}^{n}/{\mathbb{Z}}^{n}$ be a neighborhood of 0 and $\theta\in{\mathbb{T}}^{n}$. Consider the sequence $k_{1}\leq k_{2}\leq k_{3}\leq\ldots$ of all positive integers such that $k_{l}\theta\in U$. Clearly, $k_{l}\to\infty$. Moreover, it is well known that the orbit $\{k\theta\mid k\in{\mathbb{N}}\}$ is equidistributed in the closure $G$ of $\theta{\mathbb{Z}}$ in ${\mathbb{T}}^{n}$. In particular, the density of this sequence in $\mathbb{N}$} is equal to the $\dim G$-dimensional volume of $U\cap G$.
@@
The strategy of the proof is as follows. After settling some preliminaries in Sections 5.2.1 and 5.2.2, we show in Section 5.2.3 that for any element $\Phi:=\Phi_{ij}$ assertions (IR1)(IR5) are formal consequences of the requirement that $k=k_{ijl}$ satisfies a certain system of inequalities completely determined by $\Phi$. Here $d:=d_{il}$ is determined by $k$ and, of course, $\Phi$ by the first part of (IR1). Furthermore, (IR5) can be taken as the defining condition for $C_{l}$. (The existence of an infinite bounded gap sequence of solutions $k$ of this system of inequalities is a consequence of Lemma 5.1.)
@@
5.2.1. Isospectral families
In the proof we will use a general fact from symplectic linear algebra, which allows one to simplify certain index calculations. Recall that a linear map is said to be semisimple if for every eigenvalue its algebraic multiplicity is equal to the geometric multiplicity. A family of linear maps $P_{t}$, $t\in[0,1]$, is said to be isospectral if the eigenvalues of $P_{t}$ are constant. (Then, as a consequence, they are constant as a multiset.)
For such a family, the mean index $\operatorname{\hat{\mu}}(P_{t})$ and, when $P$ is non-degenerate, the Conley–Zehnder index $\mu(P_{t})$ are constant as follows immediately from the definitions; see, e.g., [Lo02, SZ]. (The non-degeneracy condition is essential: more sensitive invariants such as $\mu_{\pm}$ or $\beta_{\pm}$ need not be constant.)
Lemma 5.3.
Every $P\in\operatorname{Sp}(2m)$ can be connected in $\operatorname{Sp}(2m)$ to a semisimple linear map by an isospectral family.
As a consequence, in many index calculations we can assume that the map is semisimple by replacing $P_{0}$ by $P_{1}$. Lemma 5.3 has certainly been known in some contexts, but we are not aware of any reference and we include a proof for the sake of completeness. In fact, we need the lemma only when $P$ is elliptic and non-degenerate, but the general case might be of independent interest and the proof under these conditions is not much shorter.
Proof.
By decomposing $P$ as a direct sum depending on the type of eigenvalues, we only need to prove the lemma when $P$ is one of the following three types:
  • All eigenvalues of $P$ are elliptic (i.e., unit) and different from 1-1.
  • The only eigenvalue is 1-1.
  • All eigenvalues are hyperbolic, i.e., not unit.
We will start with the first case. Since $P$ has no real negative eigenvalues, it lies in the image of the exponential map $\exp\colon{\mathfrak{sp}}(2m)\to\operatorname{Sp}(2m)$ by [CLW, Sect. 2.6]. Thus $P=e^{A}$, where $A$ is a linear Hamiltonian operator, aka a linear Hamiltonian vector field, and it is sufficient to connect $A$ to a semisimple operator in ${\mathfrak{sp}}(2m)$. When we order the linear Darboux coordinates as $(p_{1},\ldots,p_{m},q_{1},\ldots,q_{m})$, the operator $A$ takes the form
$A=\left[\begin{array}[]{cc}A_{11}&A_{12}\\ A_{21}&A_{22}\end{array}\right],$
where $A_{11}\in{\mathfrak{gl}}(n)$ is an arbitrary matrix, $A_{22}=-A_{11}^{\top}$, and $A_{12}$ and $A_{21}$ are symmetric: $A_{12}=A_{12}^{\top}$ and $A_{21}=A_{21}^{\top}$. Furthermore, since $P$ is elliptic, $A$ has pure imaginary eigenvalues. Let us denote the spectrum of $A$ by $\sigma(A)$ with eigenvalues taken without multiplicity. Without loss of generality, we may assume that $A$ does not have semisimple eigenvalues. Then, depending on $\sigma(A)$, the operator $A$ breaks down into a direct sum of the following four normal forms:
  • (NF1)
    In this case $\sigma(A)=\{0\}$, $m$ is odd, and $A_{22}$ is the negative Jordan block of size $m$ with eigenvalue 0, i.e.,
    $A_{22}=\left[\begin{array}[]{cccc}0&-1&&\\ &\ddots&\ddots&\\ &&\ddots&-1\\ &&&0\\ \end{array}\right]$
    with the remaining empty spots standing for zeros, and $A_{12}=0=A_{21}$.
  • (NF2)
    As above, $\sigma(A)=\{0\}$ and $A_{22}$ is the negative Jordan block of size $m$ with eigenvalue 0 and $A_{12}=0$, but $m$ is arbitrary and all entries of $A_{21}$ are zero except the bottom right corner which is $\pm 1$.
  • (NF3)
    Here $\sigma(A)=\{\sqrt{-1}b\}$ with $b>0$, $m$ is even, and
    $A_{22}=\left[\begin{array}[]{cccc}B&-\tilde{I}&&\\ &\ddots&\ddots&\\ &&\ddots&-\tilde{I}\\ &&&B\\ \end{array}\right],$
    where
    $B=\left[\begin{array}[]{cc}0&-b\\ b&0\end{array}\right]\textrm{ and }\tilde{I}=\left[\begin{array}[]{cc}0&1\\ 1&0\end{array}\right]$
    and the empty spots are again zeros. Furthermore, $A_{12}=0$ and all entries of $A_{21}$ are zero except the $2\times 2$ block $\pm\tilde{I}$ at the bottom right corner.
  • (NF4)
    In this case $\sigma(A)=\{\sqrt{-1}b\}$ with $b\neq 0$, $m$ is odd, and $A_{22}$ is the negative Jordan block as in (NF1). However, $A_{12}$ and $A_{21}$ are anti-diagonal matrices:
    $A_{12}=\left[\begin{array}[]{ccccc}&&&&-b\\ &&&b&\\ &&\reflectbox{$\ddots$}&&\\ &b&&&\\ -b&&&&\\ \end{array}\right]=-A_{21}.$
These normal forms are taken from [CLW, Sect. 2.6] and are simpler to use for our purposes than their original variant from [Wi36].
Next, let us replace $\pm 1$ (but not $\pm b$) by $\pm t$ in all occurrences in these matrices. For instance, in (NF3) $\tilde{I}$ becomes $t\tilde{I}$ in $A_{22}$, etc. In each case, we obtain a family of Hamiltonian operators $A(t)$, $t\in[0,1]$, with $A(1)=A$.
It is easy to see that $A(0)$ is semisimple in all four cases. Indeed, $A(0)=0$ in (NF1) and (NF2). In (NF3) and (NF4), $A(0)$ is equivalent via a linear symplectic change of variables to a direct sum of infinitesimal rotations with eigenvalues $\pm\sqrt{-1}b$ in $m$ symplectic planes.
We claim that the four families $A(t)$ are isospectral. Let $\rho_{t}(z)=\det\big(zI-A(t)\big)$ be the characteristic polynomial. It is sufficient to show that $\rho_{t}(z)$ is independent of $t$ and this can be done by a direct calculation. In (NF1) or (NF2), we have the determinant of a block-diagonal or block-triangular matrix with triangular matrices on the diagonal, and $\rho_{t}(z)=z^{2m}$ for all $t$. Likewise, in (NF3) we have the determinant of a block-triangular matrix with block-triangular matrices on the diagonal, and $\rho_{t}(z)=(z^{2}+b^{2})^{m}$. To calculate $\rho_{t}(z)$ in (NF4), we expand the determinant in the first row and then expand each of the resulting two non-zero terms in the last row. Then $\rho_{t}(z)$ factors as the product of $z^{2}+b^{2}$ and the characteristic polynomial of a $(2m-2)\times(2m-2)$ matrix of exactly the same form, but with $b$ replaced by b-b. (The resulting matrix is not a normal form because m1m-1 is even. However, this does not matter since we are just calculating the characteristic polynomial.) Proceeding inductively, we again see that $\rho_{t}(z)=(z^{2}+b^{2})^{m}$. This concludes the proof for an elliptic map $P$ without eigenvalue 1-1.
The second case is where 1-1 is the only eigenvalue of $P$. Then P-P is covered by the previous case. Thus there exists an isospectral family connecting P-P to a semisimple symplectic map. (In fact, this map is $I$ because $1$ is the only eigenvalue of P-P.) Multiplying this family by I-I, we obtain the required isospectral family for $P$.
The remaining case is where $P$ is hyperbolic. This case can be dealt with in a fashion similar to the first two, but a much simpler proof is available. Namely, ${\mathbb{R}}^{2m}=W_{-}\oplus W_{+}$ be the decomposition of ${\mathbb{R}}^{2m}$ into subspaces corresponding to the eigenvalues $\lambda$ of $P$ with $|\lambda|<1$ and, respectively, $|\lambda|>1$. These are $P$-invariant Lagrangian subspaces. As a consequence, the symplectic form gives rise to an isomorphism $W_{+}=W_{-}^{*}$ and $P=(P_{-},\,P_{+})$ with $P_{+}=(P_{-}^{-1})^{\top}$ in this decomposition. (In other words, identifying ${\mathbb{R}}^{2m}=W_{-}\oplus W_{-}^{*}$ with $T^{*}W_{-}$ we identify $P\colon T^{*}W_{-}\to T^{*}W_{-}$ with the map induced by $P_{-}^{-1}\colon W_{-}\to W_{-}$.) Let us connect $P_{-}$ to an semisimple linear map by a family $P_{-}(t)\colon W_{-}\to W_{-}$ of isospectral linear maps. The existence of such a family immediately follows from the Jordan normal form theorem. Then $P_{t}:=\big(P_{-}(t),\,(P_{-}^{-1}(t))^{\top}\big)$ is a family of isospectral maps in $\operatorname{Sp}(2m)$ connecting $P$ to a semisimple map. This completes the proof of Lemma 5.3. ∎
Remark 5.4.
Throughout the proof of the lemma, one could have used the normal forms for symplectic matrices from [Wi36], see also [Lo02], but these normal forms have a complicated format and the proof would be more involved. The last step of the proof can also be used to directly obtain the normal forms of hyperbolic symplectic or Hamiltonian operators from the Jordan normal form theorem.
5.2.2. Preliminary observations
@@
The claim that the sequences $C_{l}$, $d_{il}$ and $k_{ijl}$ can be taken strictly increasing is a formal consequence of the rest of the theorem. Indeed, observe that every bounded gap sequence $d_{il}\to\infty$ has a bounded gap strictly increasing refinement $s\mapsto d_{il_{s}}$. Moreover, for any $c>0$ we may ensure by further thinning the sequence that $d_{il_{s+1}}-d_{l_{s}}>c$. Hence, by passing to subsequences of $d_{il}$ and $k_{ijl}$, we can guarantee that $d_{il}$ is strictly increasing, $d_{il+1}-d_{il}>c$ and these sequences still have bounded gap. Then, since $\operatorname{\hat{\mu}}\big(\Phi^{k_{ijl}}_{ij}\big)=k_{ijl}\operatorname{\hat{\mu}}(\Phi_{ij})$ and $\operatorname{\hat{\mu}}(\Phi_{ij})>0$, we conclude from (IR1) that the sequences $k_{ijl}$ are automatically strictly increasing when $c$ is sufficiently large. Likewise, the sequence $C_{l}$ is forced to be strictly increasing by (IR5) once $\eta>0$ is sufficiently small.
@@
Let us decompose $\Phi$ as the product of a loop $\varphi$ and a direct sum $\Phi_{\scriptscriptstyle{deg}}\oplus\Phi_{h}\oplus\Phi_{e}$ with the following properties. The end-point $\Phi_{\scriptscriptstyle{deg}}(1)$ is totally degenerate and $\operatorname{\hat{\mu}}(\Phi_{\scriptscriptstyle{deg}})=0$. The path $\Phi_{h}$ has a hyperbolic end-point $\Phi_{h}(1)$, while $\Phi_{e}$ is the elliptic part of $\Phi$, i.e., all eigenvalues of $\Phi_{e}(1)$ are on the unit circle and different from 1 due to the presence of $\Phi_{\scriptscriptstyle{deg}}$. We write the eigenvalues of $\Phi_{e}(1)$ as $\exp(\pm 2\pi\sqrt{-1}\lambda_{q})$, where $0<|\lambda_{q}|\leq 1/2$ and $q=1,\ldots,m^{\prime}$. Here for each pair of complex conjugate eigenvalues it is convenient to pick the sign of $\lambda_{q}$ so that the resulting eigenvalue is of the first Krein type; see [SZ]. (We will elaborate on this choice later.) Furthermore, we then require that
@@
where $\|\cdot\|$ stands for the distance to the nearest integer. We will require $\epsilon>0$ (and $\eta>0$) to be so small that
@@
Before proving the lemma, we note that (5.5) has a bounded gap sequence of solutions $k_{l}\to\infty$ by Lemma 5.1 and then, as we have already pointed out, the corresponding sequence $d=d_{l}\to\infty$ has bounded gap automatically. For any $a:=a_{ij}>0$ to find a bounded gap sequence $C_{l}\to\infty$ meeting condition (IR5), we can just pick any $C_{l}\not\in a{\mathbb{N}}$ in the range $(k_{l}a,k_{l}a+\eta)$ assuming that $\eta<a$. Furthermore, to make sure that $k$ is divisible by $D$ or any fixed integer $N>0$, it suffices to replace $\epsilon$ by $\epsilon/N$ in the definition of $U$. Then, after applying Lemma 5.1 for this new neighborhood $U$, we replace sequence of solutions $k$ by $Nk$ and have a bounded gap sequence of solutions satisfying (5.5) and divisible by $N$. As the next step, to meet the requirement of (5.6) we have $d$ replaced by $Nd$, and hence also divisible by $N$, provided that $\epsilon>0$ is sufficiently small. Thus, to complete this step of the proof, it remains to prove Lemma 5.5. The proof has a substantial overlap with [GG20, Sec. 5.2.1] and is implicitly contained in that section.
@@
In spite of a minor change of notations, these are exactly conditions (IR1)(IR3) stated for one map $\Phi=\Phi_{ij}$ with $k=k_{ijl}$, etc.
@@
$d=\sum_{q}\mathit{sign}(\lambda_{q})\big[2k|\lambda_{q}|\big],$ (5.8)
where $\cdot\,$] stands for the nearest integer. Then
@@
We will use this observation to reduce the proof to several particular cases. Throughout the argument we will assume without specifically stating that $k\in{\mathbb{N}}$ and $\ell\in{\mathbb{N}}$ and $k>\ell$ in (IR3”).
@@
By Lemma 5.3, without loss of generality, we may assume that map $\Phi(1)$ is semisimple. Indeed, consider an isospectral (i.e., with constant eigenvalues) family $\Psi_{s}$, $s\in[0,1]$, connecting $\Phi(1)$ to a semisimple linear symplectic map with the same eigenvalues as $\Phi(1)$. Let $\tilde{\Phi}_{s}\in\widetilde{\operatorname{Sp}}(2m)$ be obtained by concatenating $\Phi$ and $\Psi$. Then $\mu\big(\tilde{\Phi}_{s}^{k}\big)=\mu\big(\Phi^{k}\big)$ and $\operatorname{\hat{\mu}}\big(\tilde{\Phi}_{s}^{k}\big)=\operatorname{\hat{\mu}}\big(\Phi_{s}^{k}\big)$ for all $k$ and $s$, and hence in the proof of the theorem we can replace $\Phi$ by $\tilde{\Phi}$. We keep the notation $\Phi$.
@@
where $0<|\lambda_{q}|\leq 1/2$, $q=1,\ldots,m^{\prime}$, and $t$ ranges from 0 to $1$. Since no eigenvalue is a root of unity all $\lambda_{q}$ are irrational. Furthermore, $\operatorname{\hat{\mu}}(R_{q})=2\lambda_{q}$ and $\mu(R_{q})=\mathit{sign}(\lambda_{q})$ and $\operatorname{\hat{\mu}}(\Phi)=\sum\operatorname{\hat{\mu}}(R_{q})$. Therefore, by additivity, we only need to establish (IR2”) and (IR3”) for one rotation $R_{q}$ to complete the proof of this subcase. For the sake of brevity, set $R:=R_{q}$ and $\lambda:=\lambda_{q}$. Then these identities amount to
@@
$\operatorname{\hat{\mu}}\big(R^{k}\big)=2k\lambda\quad\textrm{and}\quad\mu\big(R^{k}\big)=\mathit{sign}(\lambda)\big(2\lfloor k|\lambda|\rfloor+1\big).$
@@
$\mathit{sign}(\lambda)\big(2\lfloor(k\pm\ell)|\lambda|\rfloor+1\big)=[2k\lambda]\pm\mathit{sign}(\lambda)\big(2\lfloor\ell|\lambda|\rfloor+1\big).$
@@
In spite of its appearance, this statement is not entirely obvious and requires a proof. An easy way to prove (5.13) is as follows. By Lemma 5.3, there exists an isospectral family $\Psi_{s}$, $s\in[0,1]$, of transformations with $\Psi_{1}=\Phi_{1}$ and $\Psi_{0}(1)$ semisimple. Then $\Psi_{0}^{k}(1)={\mathit{i}d}$, i.e., $\Psi_{0}^{k}$ is a loop, and $\operatorname{\hat{\mu}}\big(\Phi_{1}^{k}\big)=\operatorname{\hat{\mu}}\big(\Psi_{0}^{k}\big)$ because all eigenvalues of $\Psi_{s}^{k}(1)$ are equal to 1 for all $s$. Furthermore, $\Psi_{s}^{k\pm\ell}$ and $\Psi_{s}^{\ell}$ remain non-degenerate for all $s$. As a consequence, $\mu\big(\Phi_{1}^{\ell}\big)=\mu\big(\Psi_{0}^{\ell}\big)$ and $\mu\big(\Phi_{1}^{k\pm\ell}\big)=\mu\big(\Psi_{0}^{k\pm\ell}\big)$. It follows that (5.13) holds if and only if it holds for $\Psi_{0}$. The latter, however, is obvious:
@@
Let $q$ be the least common multiple of the degrees of eigenvalues of $\Phi_{0}$. Then $\Phi_{0}^{q}$ is totally degenerate. Furthermore, $q$ divides both $k$ and $\ell$, and $\Phi_{0}^{\ell}=\big(\Phi_{0}^{q}\big)^{\ell/q}$ and $\Phi_{0}^{k+\ell}=\big(\Phi_{0}^{q}\big)^{(k+\ell)/q}$ are admissible iterates of $\Phi_{0}^{q}$. Hence
$\beta_{\pm}\big(\Phi_{0}^{k+\ell}\big)=\beta_{\pm}(\Phi_{0}^{q})=\beta_{\pm}\big(\Phi_{0}^{\ell}\big)$
by (3.4) applied to $\Phi_{0}^{q}$. Therefore, using (5.13) in the third line below, we have
@@
Then a simple examination of the rest of the proof shows that (IR1”)(IR3”) still hold for $\Psi$. This completes the proof of Lemma 5.5. ∎
@@
As in Section 5.2.3, denote the elliptic eigenvalues of $\Phi_{ij}$ by $\exp(\pm 2\pi\sqrt{-1}\lambda_{ijq})$, where $0<|\lambda_{ijq}|\leq 1/2$ and $q=1,\ldots,m_{ij}^{\prime}$, and for each pair of complex conjugate eigenvalues we pick the sign of $\lambda_{ijq}$ so that the resulting eigenvalue is of the first Krein type; see [SZ]. Set $\Delta_{ij}:=\operatorname{\hat{\mu}}(\Phi_{ij})$ and let $D$ be the least common multiple of the degrees of the roots of unity among the eigenvalues of all $\Phi_{ij}$.
@@
$|k_{11}a_{11}-k_{ij}a_{ij}|$ $<\sigma\quad\textrm{ for $i+j\geq 2$},$